# Large Vision Models can solve mental rotation problems

### v1

@@ 1 Introduction
This work focuses on the mental rotation problem in the context of modern high-capacity vision models. The overview of our pipeline is illustrated in Figure 1. We investigate whether image embeddings from Google’s supervised Vision Transformer (ViT[10] and other large-scale ViT-based representation models trained with self-supervision, such as CLIP [11], DINOv2 [12], and DINOv3 [13], implicitly encode 3D structural and viewpoint information sufficient to distinguish an object from its mirrored counterpart under different rotations. In this paper, we use ViT to denote the Vision Transformer architecture in general, and ViT to refer specifically to Google’s supervised model. CLIP learns transferable image features by aligning images with text descriptions, while DINOv2 and DINOv3 are self-supervised transformers that report strong general-purpose visual features. We ask whether embeddings from these models, both at the final layer and at intermediate layers, contain information that enables mental rotation-like reasoning. Evidence of such information would suggest a form of viewpoint equivariance that preserves pose, which is appropriate for this task. By analyzing and comparing embeddings on image pairs depicting rotated or mirrored objects, we assess whether contemporary representations encode the necessary cues for mental rotation, or whether new architectures and training paradigms are required.
@@ 3 Methods
Meta’s DINOv2 [12] and DINOv3 [13] are self-supervised ViTs trained without labels on large-scale curated image datasets. DINOv2 is trained on LVD-142M, a 142M–image collection automatically curated to resemble established smaller manually curated datasets. DINOv3 scales the approach to frontier-size models with up to 7B parameters, trained on LVD-1689M, a broader 1.7B–image dataset constructed with the same principle but at a much larger scale. Both models rely on a teacher–student framework where several augmented views of the same image are generated through random resized crops and standard augmentations such as horizontal flips, color jitter, blur, and solarization [17]. The student processes all views, while the teacher, updated as an exponential moving average of the student, receives only the two global views. The teacher produces softmax distributions over a very large output space, and the student is trained with a cross-entropy loss to match these outputs, thereby enforcing augmentation-invariant representations without reconstruction. The models rely on distribution matching through image-level, patch-level, and uniformity losses. In DINOv2, this serves as the central training objective, whereas DINOv3 extends the framework with so-called Gram anchoring, a new loss term that stabilizes dense features in large-scale training [13]. This extension enables DINOv3 to retain the robust global representations of DINOv2 while further improving the stability of high-resolution dense maps.
@@ 4 Results and discussion
Across all tasks, CLIP consistently demonstrates strong performance, particularly in scenarios involving textual information. This is unsurprising given CLIP’s joint training on text–image pairs, which naturally gives it OCR-like capabilities [11]. What is especially striking, however, is that DINOv3 performs comparably well in these text-heavy conditions, despite not being explicitly trained with text supervision. In contrast, supervised ViT models perform poorly relative to their self-supervised counterparts. A possible explanation is that supervised ViTs, optimized for categorical discrimination, favor invariance over pose sensitivity, thereby discarding geometric information needed for mental-rotation tasks. For the simplest Shepard-Metzler ±0° problem, all models reach non-trivial accuracy at some point in the network. The stage at which performance peaks, however, differs: ViT Base and ViT Large exhibit the strongest representations early in the network, ViT Huge shows weak but consistent performance throughout, and the rest of the models peak at intermediate depths [2]. The difficulty of the tasks increases with increasing the relative rotational angle, in line with reaction time in human studies on mental rotation. Fruit rotations provide another example: accuracy declines as the camera angle increases, likely due to occlusion effects that disrupt consistent visual features.
The Shepard-Metzler Free condition represents the most challenging setting. Strikingly, only DINOv3 Huge shows any ability to solve this task, and even then, only at layer 18. A recurring pattern across models is the emergence of a plateau in task performance in mid-level layers, followed by a decline toward the final layers. This aligns with findings that different types of concepts are encoded at different depths of the transformer hierarchy [8, 9], and echoes observations across vision, speech, and language models that intermediate representations often provide richer or more useful structure than the final layer (e.g. [18, 19, 20, 21]). In practical terms, this indicates that the final representation layer is not always the most informative for tasks requiring geometric reasoning, and that intermediate representations may better capture the relevant structure.

### v2

@@ 1 Introduction
This work focuses on the mental rotation problem in the context of modern high-capacity vision models. The overview of our pipeline is illustrated in Figure 1. We investigate whether image embeddings from Google’s supervised Vision Transformer (ViT[8] and other large-scale ViT-based representation models trained with self-supervision, such as CLIP [16], DINOv2 [15], and DINOv3 [siméoni2025dinov3], implicitly encode 3D structural and viewpoint information sufficient to distinguish an object from its mirrored counterpart under different rotations. In this paper, we use ViT to denote the Vision Transformer architecture in general, and ViT to refer specifically to Google’s supervised model. CLIP learns transferable image features by aligning images with text descriptions, while DINOv2 and DINOv3 are self-supervised transformers that report strong general-purpose visual features. We ask whether embeddings from these models, both at the final layer and at intermediate layers, contain information that enables mental rotation-like reasoning. Evidence of such information would suggest a form of viewpoint equivariance that preserves pose, which is appropriate for this task. By analyzing and comparing embeddings on image pairs depicting rotated or mirrored objects, we assess whether contemporary representations encode the necessary cues for mental rotation, or whether new architectures and training paradigms are required.
@@ 3 Methods
Meta’s DINOv2 [15] and DINOv3 [siméoni2025dinov3] are self-supervised ViTs trained without labels on large-scale curated image datasets. DINOv2 is trained on LVD-142M, a 142M–image collection automatically curated to resemble established smaller manually curated datasets. DINOv3 scales the approach to frontier-size models with up to 7B parameters, trained on LVD-1689M, a broader 1.7B–image dataset constructed with the same principle but at a much larger scale. Both models rely on a teacher–student framework where several augmented views of the same image are generated through random resized crops and standard augmentations such as horizontal flips, color jitter, blur, and solarization [14]. The student processes all views, while the teacher, updated as an exponential moving average of the student, receives only the two global views. The teacher produces softmax distributions over a very large output space, and the student is trained with a cross-entropy loss to match these outputs, thereby enforcing augmentation-invariant representations without reconstruction. The models rely on distribution matching through image-level, patch-level, and uniformity losses. In DINOv2, this serves as the central training objective, whereas DINOv3 extends the framework with so-called Gram anchoring, a new loss term that stabilizes dense features in large-scale training [siméoni2025dinov3]. This extension enables DINOv3 to retain the robust global representations of DINOv2 while further improving the stability of high-resolution dense maps.
@@ 4 Results and discussion
Across all tasks, CLIP consistently demonstrates strong performance, particularly in scenarios involving textual information. This is unsurprising given CLIP’s joint training on text–image pairs, which naturally gives it OCR-like capabilities [16]. What is especially striking, however, is that DINOv3 performs comparably well in these text-heavy conditions, despite not being explicitly trained with text supervision. In contrast, supervised ViT models perform poorly relative to their self-supervised counterparts. A possible explanation is that supervised ViTs, optimized for categorical discrimination, favor invariance over pose sensitivity, thereby discarding geometric information needed for mental-rotation tasks. For the simplest Shepard-Metzler ±0° problem, all models reach non-trivial accuracy at some point in the network. The stage at which performance peaks, however, differs: ViT Base and ViT Large exhibit the strongest representations early in the network, ViT Huge shows weak but consistent performance throughout, and the rest of the models peak at intermediate depths. The difficulty of the tasks increases with increasing the relative rotational angle. This parallels human reaction times in mental rotation [17], and is consistent with findings that humans and deep networks face similar difficulties with geometric transformations [12]. Fruit rotations show a similar pattern, with accuracy declining as object occlusion increases.
The Shepard-Metzler Free condition represents the most challenging setting. Strikingly, only DINOv3 Huge shows any ability to solve this task, and even then, only at layer 18. A recurring pattern across models is the emergence of a plateau in task performance in mid-level layers, followed by a decline toward the final layers. This aligns with findings that different types of concepts are encoded at different depths of the transformer hierarchy [6, 21], and echoes observations across vision, speech, and language models that intermediate representations often provide richer or more useful structure than the final layer (e.g. [20, 7, 18]). In practical terms, this indicates that the final representation layer is not always the most informative for tasks requiring geometric reasoning, and that intermediate representations may better capture the relevant structure.